Please help me solve this problem.



The probabilities of ordering 0,1,2,3, or 4 toppings on a pizza are 2/10, 5/10, 2/10, 7/100, 3/100 respectively. What is the mathematical expectation for the number of toppings ordered on a pizza? (State the answer in integer/decimal form, i.e. 1.30, .35, etc.)

Answer by stanbon(75887)

The probabilities of ordering 0,1,2,3, or 4 toppings on a pizza are 2/10, 5/10, 2/10, 7/100, 3/100 respectively. What is the mathematical expectation for the number of toppings ordered on a pizza? (State the answer in integer/decimal form, i.e. 1.30, .35, etc.)
----
E(x) = (20*0+50*1+20*2+7*3+3*4)/100 = 123/100 = 1.23
=========================
Cheers,
Stan H.